OpenStudy (anonymous):

LagrangeSon678 yes i wrote the problem correctly it tells you to use rational exponents to write the problem i gave as a radical expression and then choose from the following four answers

OpenStudy (anonymous):

thats if i am understanding this problem correctly

OpenStudy (anonymous):

that is not one of the choices the others are:\[\sqrt[4]{175616}\]\[\sqrt[8]{56}\]\[\sqrt[4]{392}\]\[1/\sqrt{392^{4}}\]

OpenStudy (anonymous):

you can add the exponents unless the bases are the same.

OpenStudy (anonymous):

cant*

OpenStudy (anonymous):

since 7 and 8 are different you cant add them like that.

OpenStudy (anonymous):

\[8^\frac{1}{4}*7^\frac{1}{2} = 8^\frac{1}{4}*7^\frac{2}{4} = 8^\frac{1}{4}*49^\frac{1}{4} = (8*49)^\frac{1}{4} = 392^\frac{1}{4}\]

OpenStudy (anonymous):

joemath314159 the correct ans. is: 1/392^4

OpenStudy (anonymous):

where is the original problem? i was going off what someone posted.

OpenStudy (anonymous):

Use rational exponent to write \[\sqrt[4]{8} * \sqrt[2]{7}\] as a radical expression and choose from the following answers I posted ahead and above

myininaya (myininaya):

joe i just did that lol on his previous post

myininaya (myininaya):

lol u won :(

OpenStudy (anonymous):

i dont feel like i won lol >.> theres nothing wrong with my answer.

myininaya (myininaya):

speed racer

OpenStudy (anonymous):

\[392^\frac{1}{4} = \sqrt[4]{392}\]

myininaya (myininaya):

yours say 6 min ago my said 2 min ago so i lost

OpenStudy (anonymous):

so guys what the finaal ans.

OpenStudy (anonymous):

you still get a medal!

myininaya (myininaya):

its the third choice as i said on the previous post and as joe said here

OpenStudy (anonymous):

which is 4√392

OpenStudy (anonymous):

right

OpenStudy (anonymous):

thnks have a good night that was my last problem for a quiz

OpenStudy (anonymous):

night :)

OpenStudy (anonymous):

im just waking up >.>

OpenStudy (anonymous):

ok well good morining thnks again